# Train for the Worst, Plan for the Best: Understanding Token Ordering in Masked Diffusions

### v2

@@ 4.2 Adaptive MDM inference
MDM (Top-$K$ probability) 98.5 %
MDM (Top-$K$ prob. margin) 98.3 %
4.3 Eliciting sequence-dependent reasoning paths using adaptive MDM inference
@@
We compare ARMs and MDMs for Sudoku in Table 2 and Zebra puzzles in Table 3.222A prior work (Ye et al., 2024) reported that a $6$M MDM with Top-$K$ inference achieves 100% accuracy on Sudoku. Given that a 6M MDM with Top-$K$ only achieves 18.51% on our dataset (Table 2), this suggests that the Sudoku dataset in (Ye et al., 2024) is significantly easier than ours. We observe that for both, Top-$K$ probability margin-based adaptive MDM inference not only outperforms the ARM trained without ordering information, but it even outperforms the ARM trained with ordering information! This shows that the unsupervised way of finding the correct order and solving such logic puzzles using adaptive MDM inference outperforms the supervised way of finding the correct order and solving such puzzles using an ARM, and is significantly less computationally intensive.
@@ 4.4 Easy to hard generalization
Method ##\##Param Accuracy
@@
MDM (Top-$K$ probability) 9.44 %
MDM (Top-$K$ prob. margin) 49.88 %
We measure the accuracy of MDMs and ARMs on the hard test set and present the results in Table 4. We see that the Top-$K$ probability margin-based adaptive MDM inference strategy (49.88%) again significantly outperforms ARMs trained with order information (32.57%). In particular, although the accuracy drops for both methods due to the more challenging test set, MDMs with adaptive inference appear to be more robust to this distribution shift than ARMs. We believe this is due to the fact that MDMs try to solve a significantly higher number of infilling problems than ARMs ($\exp(L)$ compared to $L$) and therefore are able to extract knowledge about the problem more efficiently than ARMs.
@@ 5 Conclusion
In this work, we examined the impact of token ordering on training and inference in MDMs. We provided theoretical and experimental evidence that MDMs train on hard masking problems. We also demonstrated that adaptive inference strategies can be used to sidestep these hard problems. For logic puzzles, we find that this leads to dramatic improvements in performance not just over vanilla MDMs, but even over ARMs trained with teacher forcing to learn the right order of decoding.
An important direction for future work is to explore settings beyond logic puzzles where adaptive inference can help MDMs match or surpass ARMs. For these, it may be crucial to go beyond the relatively simple adaptive strategies like Top-$K$ and Top-$K$ probability margin considered here.
@@ Acknowledgements.
JK thanks Kiwhan Song for discussions about MDM training. KS and VK are supported by the NSF AI Institute for Foundations of Machine Learning (IFML). KS thanks Nishanth Dikkala for the initial discussions about the project. SC is supported by the Harvard Dean’s Competitive Fund for Promising Scholarship and thanks Brice Huang and Sidhanth Mohanty for enlightening discussions about computational-statistical tradeoffs for planted CSPs.
@@ Appendix B Technical details from Section 3
Throughout this section, we use $x^{i}$ to denote the $i$-th coordinate of the vector $x$ and $z{(j)}$ to denote the $j$-th example. The $i$-th coordinate of the vector $z{(j)}$ is denoted by $z{(j)}^{i}$.
@@ B.1 Additional example: sparse parity observations
Let $10<\alpha<$ 1 be an arbitrary absolute constant, and let $\eta=1/\mathrm{poly}(N)$ be sufficiently large. Let $x$ be a sample from a L&O distribution $p_{\rm data}$ with noisy parity observations as defined in Example B.1. Suppose each token is independently masked with probability $\alpha$, and $M$ is the set of indices for the masked tokens. If $1-1/N\leq\alpha\leq 1-1/2N$, then under the Sparse Learning Parity with Noise (SLPN) assumption (see Definition B.3), with constant probability over $M$, no polynomial-time algorithm can solve the resulting masking problem of predicting any of the masked tokens among $x^{\pi(1)},\ldots,x^{\pi(N)}$ given $x[M]$.
@@
Given input dimension $N$, noise parameter $20<\eta<1/$ 2, and sample size $P$, an instance of the Sparse Learning Parity with Noise (SLPN) problem is generated as follows:
@@
  • We observe $P$ examples of the form $(x{(i)},y{(i)})$ where $x{(i)}$ is sampled independently and uniformly at random from $k$-sparse bitstrings in $\{0,1\}^{N}$, and $y$ is given by $\epsilon_{i}+\langle x{(i)},x\rangle\pmod{2}$, where $\epsilon_{i}$ is $1$ with probability $\eta$ and 00 otherwise.
@@
The SLPN assumption is that for any $P=N^{(1-\rho)k/2}$ for constant $10<\rho<$ 1, and any sufficiently large inverse polynomial noise rate $\eta$, no $\mathrm{poly}(N)$-time algorithm can recover $x$ with high probability.
@@
With probability at least $1-(1-1/N)^{N}\geq\Omega(1)$, all of the variable tokens $x^{\pi(i)}$ for $i\leq N$ are masked. Independently, the number of unmasked tokens among the observation tokens $\mathcal{O}_{j}$ is distributed as $\mathrm{Bin}(P,1-\alpha)$, so by a Chernoff bound, with probability at least $1-e^{-\Omega(P/N^{2})}=1-1/\mathrm{poly}(N)$ we have that at least $P/4N=\Omega(N\log N)$ observation tokens are unmasked. The masking problem in this case amounts to an instance of SLPN with input dimension $N$ and sample size in $[\Omega(N\log N),O(N^{0.49k})]$. Because of the lower bound on the sample size, prediction of $\mathbf{x}^{M}$ is information-theoretically possible. Because of the upper bound on the sample size, the SLPN assumption makes it computationally hard. As a result, estimating the posterior mean on any entry of $\mathbf{x}^{M}$ given the unmasked tokens is computationally hard as claimed. ∎
@@ B.2 Additional example: random slab observations
Let $m=2$ and $P=\gamma N^{2}$ for constant $\gamma>0$. Fix slab width $\beta$ and vectors $z{(1)},\ldots,z{(P)}$ sampled independently from $\mathcal{N}(0,I)$. For each $j\in[P]$, define the corresponding observation $\mathcal{O}_{j}(x)$ to be deterministically $1$ if $|\langle z{(j)},2x-\mathbf{1}\rangle|\leq\beta\sqrt{N}$, and deterministically 00 otherwise.
@@
Let $p_{\rm data}$ be a L&O distribution with random slab observations as defined in Example B.4, with parameter $\gamma>0$ and slab width $\beta>0$. There exists a constant $c_{\beta}>0$ such that for any absolute constant $0<c<c_{\beta}$, if $1-c_{\beta}N/2P\leq\alpha\leq 1-cN/P$ and $\gamma>c_{\beta}$, the following holds. Let $p^{\prime}_{\rm data}$ denote the distribution given by independently masking every coordinate in $p_{\rm data}$ with probability $\alpha$. Then any $(1-\tilde{\Omega}(1/\sqrt{N}))$-stable algorithm, even one not based on masked diffusion, which takes as input a sample $x^{\prime}$ from $p^{\prime}_{\rm data}$ and, with probability $1-o(1)$ outputs a Wasserstein-approximate [^3] Here the notion of approximation is $o(1)$-closeness in Wasserstein-2 distance. sample from $p_{\rm data}$ conditioned on the unmasked tokens in $x^{\prime}$, must run in super-polynomial time.
@@
  • Nature samples $\sigma$ uniformly at random from {±$\{\pm 1\}^{N}$
  • For each $j=1,\ldots,P=\lfloor\alpha N\rfloor$, we sample $z{(j)}$ independently from $\mathcal{N}(0,I_{N})$ conditioned on satisfying $|\langle z{(j)},\sigma\rangle|\leq\beta\sqrt{N}$.
@@
Given a matrix $Z\sim\mathcal{N}(0,1)^{\otimes P\times N}$, define $Z_{t}=tZ+\sqrt{1-t^{2}}Z^{\prime}$ for independent $Z^{\prime}\sim\mathcal{N}(0,1)^{\otimes P\times N}$. A randomized algorithm $\mathcal{A}$ which takes as input $Z\in\mathbb{R}^{P\times N}$ and outputs an element of {±$\{\pm 1\}^{N}$ is said to be $t_{N}$-stable if $\lim_{N\to\infty}W_{2}(\mathrm{law}(\mathcal{A}(Z)),\mathrm{law}(\mathcal{A}(Z_{t})))=0$.
@@
For any constant $\beta>0$, there exists $c_{\beta}>0$ such that the following holds for all constants $0<\alpha<c_{\beta}$. For $t_{N}\leq 1-\Omega(\log^{2}(n)/n^{2})$, any $t_{N}$-stable randomized algorithm $\mathcal{A}$ which takes as input $Z=(z{(1)},\ldots,z{(P)})$ and outputs an element of {±$\{\pm 1\}^{N}$ will fail to sample from the posterior on $\sigma$ conditioned on $Z$ in the symmetric Ising perceptron model to Wasserstein error $o(\sqrt{N})$.
@@
By a union bound, with probability at least $1-(1-\alpha)N\geq 1-c_{\beta}N^{2}/P\geq 1-c_{\beta}/\gamma$ over a draw $x^{\prime}\sim p^{\prime}_{\rm data}$, all of the $x^{\pi(i)}$ tokens are masked. The number of unmasked tokens in $x^{\prime}$ among the observations $\mathcal{O}_{j}$ is distributed as $\mathrm{Bin}(P,1-\alpha)$. By a Chernoff bound, this is in $[3cN/4,3c_{\beta}N/4]$ with at least constant probability. The claim then follows immediately from Theorem B.8 above. ∎
B.3 Proof of Proposition 3.3: sparse predicate observations
@@
$\textrm{MS}^{S\to i}_{c}[t+1]$ $\propto\sum_{\overline{\sigma}\in\{1,\ldots,m\}^{S\backslash i}}g(\overline{\sigma}\cup_{i}c)\prod_{j:i\neq j\in S}\textrm{MS}^{j\to S}_{\overline{\sigma}_{j}}[t]\,,$ (5)
@@
The quantity $\sum_{\overline{\sigma}\in\{1,\ldots,m\}^{[k]}\backslash i}g(\overline{\sigma}\cup_{i}c)$ is constant across all $c\in\{1,\ldots,m\}$ and $i\in[k]$.
@@ C.2.1 Experiment on L&O-NAE-SAT distribution
We consider the L&O-NAE-SAT distribution with $(N,P)=(20,280)$. For each example sequence from L&O-NAE-SAT, we pad the last $212$ tokens with an additional token value of $2$. We employ a $19$M MDM with RoPE and a maximum sequence length of $512$. Then, this MDM is trained for $2\times 10^{3}$ iterations. To attain a proxy MDM for the Bayes optimal predictor, we further train it for $5\times 10^{4}$ iterations.
To measure the error across different tasks, we consider the following setup. For each $\ell\in[1,N-1]$, we randomly mask $\ell$ tokens in the latent positions and ×$\ell\times(P/N)$ tokens in the observed positions. Across all masked prediction positions, $\ell(1+P/N)$, we measure the error for each position. For certainty, we repeat this process $1000$ times. The result in Figure 2 corresponds to the case when $\ell=11$, and we observe the same tendency for other values of $\ell$.
@@ C.2.2 Experiment on text data
We take a $170$M MDM pretrained with text data for a baseline model. To measure the performance imbalance between likelihood modeling tasks
@@
As done in the experiments in Section 3.2, we sample $\pi$s from three different distributions: $\mathrm{Unif}(\mathbb{S}_{L})$, (Closer), the point mass of identical distribution. For each case, we calculate the expectation over $1024$ samples of $x_{0}\sim p_{\rm{data}}$.

### v3

@@ 4.2 Adaptive MDM inference
MDM (Top probability) 98.5 %
MDM (Top prob. margin) 98.3 %
4.3 Eliciting sequence-dependent reasoning paths using adaptive MDM inference in logic puzzles
@@
We compare ARMs and MDMs for Sudoku in Table 2 and Zebra puzzles in Table 3. We observe that for both, Top probability margin-based adaptive MDM inference not only outperforms the ARM trained without ordering information, but it even outperforms the ARM trained with ordering information! This shows that the unsupervised way of finding the correct order and solving such logic puzzles using adaptive MDM inference outperforms the supervised way of finding the correct order and solving such puzzles using an ARM, and is significantly less computationally intensive.
4.4 Adaptive MDM inference on natural language tasks
To examine the effect of different inference strategies on text benchmarks, we adapted LLaDA, the 8B MDM model from (Nie et al., 2025). We compare three inference strategies: vanilla, top probability, and top probability margin. The results are presented in Table 4.
We see that both adaptive MDM inference strategies, top probability and top probability margin, consistently outperform vanilla MDM inference. Notably, top probability margin demonstrates a clear advantage over top probability in challenging tasks like HumanEval-Multiline (infill), HumanEval-Split Line (infill), and Math. This is because Top probability margin provides a more reliable estimate of uncertainty when multiple tokens have similar probabilities, a frequent occurrence in these difficult tasks. These results further underscore the potential for developing new, sophisticated adaptive inference strategies for various tasks. We provide experimental details in Appendix D.3.
Table 4: Performance of different inference strategies for LLaDa 8B model on coding and math tasks.
Method HumanEval-Single HumanEval-Multi HumanEval-Split Math MMLU ROCStories
Vanilla 31.8% 16.5% 14.2% 28.5% 33.2% 21.23%
Top probability 32.9% 20.8% 18.4% 31.3% 36.5% 21.10%
Top prob. margin 33.5% 25.4% 22.3% 34.3% 35.4% 21.41%
@@ 4.5 Easy to hard generalization
Method #\#Param Accuracy
@@
MDM (Top probability) 9.44 %
MDM (Top prob. margin) 49.88 %
We measure the accuracy of MDMs and ARMs on the hard test set and present the results in Table 5. We see that the Top probability margin-based adaptive MDM inference strategy (49.88%) again significantly outperforms ARMs trained with order information (32.57%). In particular, although the accuracy drops for both methods due to the more challenging test set, MDMs with adaptive inference appear to be more robust to this distribution shift than ARMs. We believe this is due to the fact that MDMs try to solve a significantly higher number of infilling problems than ARMs ($\exp(L)$ compared to $L$) and therefore are able to extract knowledge about the problem more efficiently than ARMs.
@@ 5 Conclusion
In this work, we examined the impact of token generation order on training and inference in MDMs. We provided theoretical and experimental evidence that MDMs train on hard masking problems. We also demonstrated that adaptive inference strategies can be used to sidestep these hard problems. For logic puzzles, we find that this leads to dramatic improvements in performance not just over vanilla MDMs, but even over ARMs trained with teacher forcing to learn the right order of decoding. An important direction for future work is to go beyond the relatively simple adaptive strategies to find a better generation order like top probability and top probability margin considered here.
@@ Acknowledgements.
JK thanks Kiwhan Song for discussions about MDM training. KS and VK are supported by the NSF AI Institute for Foundations of Machine Learning (IFML). KS and VK thank the computing support on the Vista GPU Cluster through the Center for Generative AI (CGAI) and the Texas Advanced Computing Center (TACC) at UT Austin. KS thanks Nishanth Dikkala for the initial discussions about the project. SK acknowledges: this work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence and support from the Office of Naval Research under award N00014-22-1-2377. SC is supported by the Harvard Dean’s Competitive Fund for Promising Scholarship and thanks Brice Huang and Sidhanth Mohanty for enlightening discussions about computational-statistical tradeoffs for planted CSPs.
@@ Appendix B Technical details from Section 3
Throughout this section, we use $x^{i}$ to denote the $i$-th coordinate of the vector $x$ and z(j)z{(j)} to denote the $j$-th example. The $i$-th coordinate of the vector z(j)z{(j)} is denoted by $z{(j)}^{i}$.
@@ B.1 Additional example: sparse parity observations
Let $0<\alpha<1$ be an arbitrary absolute constant, and let $\eta=1/\mathrm{poly}(N)$ be sufficiently large. Let $x$ be a sample from a L&O distribution $p_{\rm data}$ with noisy parity observations as defined in Example B.1. Suppose each token is independently masked with probability $\alpha$, and $M$ is the set of indices for the masked tokens. If $1-1/N\leq\alpha\leq 1-1/2N$, then under the Sparse Learning Parity with Noise (SLPN) assumption (see Definition B.3), with constant probability over $M$, no polynomial-time algorithm can solve the resulting masking problem of predicting any of the masked tokens among $x^{\pi(1)},\ldots,x^{\pi(N)}$ given $x[M]$.
@@
Given input dimension $N$, noise parameter $0<\eta<1/2$, and sample size $P$, an instance of the Sparse Learning Parity with Noise (SLPN) problem is generated as follows:
@@
  • We observe $P$ examples of the form (x(i),y(i))(x{(i)},y{(i)}) where x(i)x{(i)} is sampled independently and uniformly at random from $k$-sparse bitstrings in $\{0,1\}^{N}$, and $y$ is given by $\epsilon_{i}+\langle x{(i)},x\rangle\pmod{2}$, where $\epsilon_{i}$ is $1$ with probability $\eta$ and 0 otherwise.
@@
The SLPN assumption is that for any $P=N^{(1-\rho)k/2}$ for constant $0<\rho<1$, and any sufficiently large inverse polynomial noise rate $\eta$, no $\mathrm{poly}(N)$-time algorithm can recover $x$ with high probability.
@@
With probability at least $1-(1-1/N)^{N}\geq\Omega(1)$, all of the variable tokens $x^{\pi(i)}$ for $i\leq N$ are masked. Independently, the number of unmasked tokens among the observation tokens $\mathcal{O}_{j}$ is distributed as $\mathrm{Bin}(P,1-\alpha)$, so by a Chernoff bound, with probability at least $1-e^{-\Omega(P/N^{2})}=1-1/\mathrm{poly}(N)$ we have that at least $P/4N=\Omega(N\log N)$ observation tokens are unmasked. The masking problem in this case amounts to an instance of SLPN with input dimension $N$ and sample size in $\Omega(N\log N),O(N^{0.49k})]$. Because of the lower bound on the sample size, prediction of $\mathbf{x}^{M}$ is information-theoretically possible. Because of the upper bound on the sample size, the SLPN assumption makes it computationally hard. As a result, estimating the posterior mean on any entry of $\mathbf{x}^{M}$ given the unmasked tokens is computationally hard as claimed. ∎
@@ B.2 Additional example: random slab observations
Let $m=2$ and $P=\gamma N^{2}$ for constant $\gamma>0$. Fix slab width $\beta$ and vectors $z{(1)},\ldots,z{(P)}$ sampled independently from $\mathcal{N}(0,I)$. For each $j\in[P]$, define the corresponding observation $\mathcal{O}_{j}(x)$ to be deterministically $1$ if $|\langle z{(j)},2x-\mathbf{1}\rangle|\leq\beta\sqrt{N}$, and deterministically 0 otherwise.
@@
Let $p_{\rm data}$ be a L&O distribution with random slab observations as defined in Example B.4, with parameter $\gamma>0$ and slab width $\beta>0$. There exists a constant $c_{\beta}>0$ such that for any absolute constant $0<c<c_{\beta}$, if $1-c_{\beta}N/2P\leq\alpha\leq 1-cN/P$ and $\gamma>c_{\beta}$, the following holds. Let $p^{\prime}_{\rm data}$ denote the distribution given by independently masking every coordinate in $p_{\rm data}$ with probability $\alpha$. Then any $(1-\tilde{\Omega}(1/\sqrt{N}))$-stable algorithm, even one not based on masked diffusion, which takes as input a sample $x^{\prime}$ from $p^{\prime}_{\rm data}$ and, with probability 1o(1)1-o(1) outputs a Wasserstein-approximate [^3] Here the notion of approximation is $o(1)$-closeness in Wasserstein-2 distance. sample from $p_{\rm data}$ conditioned on the unmasked tokens in $x^{\prime}$, must run in super-polynomial time.
@@
  • Nature samples $\sigma$ uniformly at random from $\{\pm 1\}^{N}$
  • For each $j=1,\ldots,P=\lfloor\alpha N\rfloor$, we sample z(j)z{(j)} independently from $\mathcal{N}(0,I_{N})$ conditioned on satisfying $|\langle z{(j)},\sigma\rangle|\leq\beta\sqrt{N}$.
@@
Given a matrix $Z\sim\mathcal{N}(0,1)^{\otimes P\times N}$, define $Z_{t}=tZ+\sqrt{1-t^{2}}Z^{\prime}$ for independent $Z^{\prime}\sim\mathcal{N}(0,1)^{\otimes P\times N}$. A randomized algorithm $\mathcal{A}$ which takes as input $Z\in\mathbb{R}^{P\times N}$ and outputs an element of $\{\pm 1\}^{N}$ is said to be $t_{N}$-stable if $\lim_{N\to\infty}W_{2}(\mathrm{law}(\mathcal{A}(Z)),\mathrm{law}(\mathcal{A}(Z_{t})))=0$.
@@
For any constant $\beta>0$, there exists $c_{\beta}>0$ such that the following holds for all constants $0<\alpha<c_{\beta}$. For $t_{N}\leq 1-\Omega(\log^{2}(n)/n^{2})$, any $t_{N}$-stable randomized algorithm $\mathcal{A}$ which takes as input $Z=(z{(1)},\ldots,z{(P)})$ and outputs an element of $\{\pm 1\}^{N}$ will fail to sample from the posterior on $\sigma$ conditioned on $Z$ in the symmetric Ising perceptron model to Wasserstein error $o(\sqrt{N})$.
@@
By a union bound, with probability at least $1-(1-\alpha)N\geq 1-c_{\beta}N^{2}/P\geq 1-c_{\beta}/\gamma$ over a draw $x^{\prime}\sim p^{\prime}_{\rm data}$, all of the $x^{\pi(i)}$ tokens are masked. The number of unmasked tokens in $x^{\prime}$ among the observations $\mathcal{O}_{j}$ is distributed as $\mathrm{Bin}(P,1-\alpha)$. By a Chernoff bound, this is in $3cN/4,3c_{\beta}N/4]$ with at least constant probability. The claim then follows immediately from Theorem B.8 above. ∎
B.3 Proof outline of Proposition 3.3
Additional examples of the hardness.
The above setup can also be generalized to capture Bayesian constraint satisfaction problems (Montanari, 2008; Liu et al., 2022), one notable example of which is the stochastic block model (Decelle et al., 2011). There are analogous predictions for the onset of hardness of inference, which can likewise be translated to hardness of masking problems for seemingly benign L&O distributions. In Appendix B.1 and B.2, we give two more examples of L&O distributions for which order-aware training is tractable yet order-agnostic training of the MDM is computationally hard.
First, we consider L&O distributions whose observations are sparse, noisy parities in the latents and deduce hardness for order-agnostic training from the Sparse Learning Parity with Noise assumption (Alekhnovich, 2003). We then consider L&O distributions whose observations are generalized linear models in the latents, and deduce hardness for a large class of efficient algorithms from existing results on Lipschitz hardness (Alaoui & Gamarnik, 2024) for the symmetric binary perceptron (Aubin et al., 2019).
B.4 Proof of Proposition 3.3: sparse predicate observations
@@
[…]
@@
[…]
@@ C.2.1 Experiment on L&O-NAE-SAT distribution
We consider the L&O-NAE-SAT distribution with $(N,P)=(20,280)$. For each example sequence from L&O-NAE-SAT, we pad the last $212$ tokens with an additional token value of $2$. We employ a 1919M MDM with RoPE and a maximum sequence length of $512$. Then, this MDM is trained for $2\times 10^{3}$ iterations. To attain a proxy MDM for the Bayes optimal predictor, we further train it for $5\times 10^{4}$ iterations.
To measure the error across different tasks, we consider the following setup. For each $\ell\in[1,N-1]$, we randomly mask $\ell$ tokens in the latent positions and $\ell\times(P/N)$ tokens in the observed positions. Across all masked prediction positions, $\ell(1+P/N)$, we measure the error for each position. For certainty, we repeat this process $1000$ times. The result in Figure 2 corresponds to the case when $\ell=11$, and we observe the same tendency for other values of $\ell$.
@@ C.2.2 Experiment on text data
We take a 170170M MDM pretrained with text data for a baseline model. To measure the performance imbalance between likelihood modeling tasks
@@
[…]